Message 5.7 Manual

by Lars Magne Ingebrigtsen

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1996 Free Software Foundation, Inc. Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions.

Message

All message composition from Gnus (both mail and news) takes place in Message mode buffers.

This manual corresponds to Message 5.7. Message is distributed with the Gnus distribution bearing the same version number as this manual has.

2 Message 5.7 Manual

1 Interface

When a program (or a person) wants to respond to a message – reply, follow up, forward, cancel – the program (or person) should just put point in the buffer where the message is and call the required command. Message will then pop up a new message mode buffer with appropriate headers filled out, and the user can edit the message before sending it.

1.1 New Mail Message

The message-mail command pops up a new message buffer.

Two optional parameters are accepted: The first will be used as the To header and the second as the Subject header. If these are nil, those two headers will be empty.

1.2 New News Message

The message-news command pops up a new message buffer.

This function accepts two optional parameters. The first will be used as the Newsgroups header and the second as the Subject header. If these are nil, those two headers will be empty.

1.3 Reply

The message-reply function pops up a message buffer that’s a reply to the message in the current buffer.

Message uses the normal methods to determine where replies are to go (see Section 5.1 [Responses], page 21), but you can change the behavior to suit your needs by fiddling with

the message-reply-to-function variable.

If you want the replies to go to the Sender instead of the From, you could do something like this:

     (setq message-reply-to-function
           (lambda ()
            (cond ((equal (mail-fetch-field "from") "somebody")
                    (mail-fetch-field "sender"))
                  (t
                   nil))))

This function will be called narrowed to the head of the article that is being replied to.

As you can see, this function should return a string if it has an opinion as to what the To header should be. If it does not, it should just return nil, and the normal methods for determining the To header will be used.

This function can also return a list. In that case, each list element should be a cons, where the car should be the name of an header (eg. Cc) and the cdr should be the header value (eg. ‘larsi@ifi.uio.no’). All these headers will be inserted into the head of the outgoing mail.

1.4 Wide Reply

The message-wide-reply pops up a message buffer that’s a wide reply to the message in the current buffer. A wide reply is a reply that goes out to all people listed in the To, From (or Reply-to) and Cc headers.

Message uses the normal methods to determine where wide replies are to go, but you can change the behavior to suit your needs by fiddling with the message-wide-reply-to-function. It is used in the same way as message-reply-to-function (see Section 1.3 [Reply], page 3).

Addresses that match the rmail-dont-reply-to-names regular expression will be removed from the Cc header.

1.5 Followup

The message-followup command pops up a message buffer that’s a followup to the message in the current buffer.

Message uses the normal methods to determine where followups are to go, but you can change the behavior to suit your needs by fiddling with the message-followup-to-function. It is used in the same way as message-reply-to-function (see Section 1.3 [Reply], page 3).

The message-use-followup-to variable says what to do about Followup-To headers. If it is use, always use the value. If it is ask (which is the default), ask whether to use the value. If it is t, use the value unless it is ‘poster’. If it is nil, don’t use the value.

1.6 Canceling News

The message-cancel-news command cancels the article in the current buffer.

1.7 Superseding

The message-supersede command pops up a message buffer that will supersede the message in the current buffer.

Headers matching the message-ignored-supersedes-headers are removed before popping up the new message buffer. The default is

‘^Path:\\|^Date\\|^NNTP-Posting-Host:\\|^Xref:\\|^Lines:\\|
^Received:\\|^X-From-Line:\\|Return-Path:\\|^Supersedes:’.

1.8 Forwarding

The message-forward command pops up a message buffer to forward the message in the current buffer. If given a prefix, forward using news.

message-forward-start-separator

Delimiter inserted before forwarded messages. The default is

          ‘------- Start of forwarded message -------\n’.

message-forward-end-separator

Delimiter inserted after forwarded messages. The default is

‘------- End of forwarded message -------\n’.
message-signature-before-forwarded-message

If this variable is t, which it is by default, your personal signature will be inserted before the forwarded message. If not, the forwarded message will be inserted first in the new mail.

message-included-forward-headers

Regexp matching header lines to be included in forwarded messages.

message-make-forward-subject-function

A list of functions that are called to generate a subject header for forwarded messages. The subject generated by the previous function is passed into each successive function. The provided functions are:

message-forward-subject-author-subject

Source of article (author or newsgroup), in brackets followed by the subject.

message-forward-subject-fwd

Subject of article with ‘Fwd:’ prepended to it.

message-wash-forwarded-subjects

If this variable is t, the subjects of forwarded messages have the evidence of previous forwards (such as ‘Fwd:’, ‘Re:’, ‘(fwd)’) removed before the new subject is constructed. The default value is nil.

1.9 Resending

The message-resend command will prompt the user for an address and resend the message in the current buffer to that address.

Headers that match the message-ignored-resent-headers regexp will be removed before sending the message. The default is ‘^Return-receipt’.

1.10 Bouncing

The message-bounce command will, if the current buffer contains a bounced mail message, pop up a message buffer stripped of the bounce information. A bounced message is typically a mail you’ve sent out that has been returned by some mailer-daemon as undeliverable.

Headers that match the message-ignored-bounced-headers regexp will be removed before popping up the buffer. The default is ‘^\\(Received\\|Return-Path\\):’.

6 Message 5.7 Manual

2 Commands

2.1 Header Commands

All these commands move to the header in question. If it doesn’t exist, it will be inserted. C-c ? Describe the message mode.

C-c C-f C-t
          Go to the To header (message-goto-to).
C-c C-f C-b
          Go to the Bcc header (message-goto-bcc).
C-c C-f C-f
          Go to the Fcc header (message-goto-fcc).
C-c C-f C-c
          Go to the Cc header (message-goto-cc).
C-c C-f C-s
          Go to the Subject header (message-goto-subject).
C-c C-f C-r
          Go to the Reply-To header (message-goto-reply-to).
C-c C-f C-n
          Go to the Newsgroups header (message-goto-newsgroups).
C-c C-f C-d
          Go to the Distribution header (message-goto-distribution).
C-c C-f C-o
          Go to the Followup-To header (message-goto-followup-to).
C-c C-f C-k
          Go to the Keywords header (message-goto-keywords).
C-c C-f C-u
          Go to the Summary header (message-goto-summary).

2.2 Movement

C-c C-b Move to the beginning of the body of the message (message-goto-body). C-c C-i Move to the signature of the message (message-goto-signature).

2.3 Insertion

C-c C-y Yank the message that’s being replied to into the message buffer (message-

yank-original).

C-c C-q Fill the yanked message (message-fill-yanked-message). Warning: Can severely mess up the yanked text if its quoting conventions are strange. You’ll quickly get a feel for when it’s safe, though. Anyway, just remember that C-x u (undo) is available and you’ll be all right.

C-c C-w Insert a signature at the end of the buffer (message-insert-signature).

message-ignored-cited-headers

All headers that match this regexp will be removed from yanked messages. The default is ‘.’, which means that all headers will be removed.

message-citation-line-function

Function called to insert the citation line. The default is message-insert-citation-line, which will lead to citation lines that look like:

     Hallvard B Furuseth <h.b.furuseth@usit.uio.no> writes:

Point will be at the beginning of the body of the message when this function is called.

message-yank-prefix

When you are replying to or following up an article, you normally want to quote the person you are answering. Inserting quoted text is done by yanking, and each quoted line you yank will have message-yank-prefix prepended to it. The default is ‘> ’. If it is nil, just indent the message.

message-indentation-spaces

Number of spaces to indent yanked messages.

message-cite-function

Function for citing an original message. The default is message-cite-original, which simply inserts the original message and prepends ‘> ’ to each

line. message-cite-original-without-signature does the same, but elides

the signature. You can also set it to sc-cite-original to use Supercite.

message-indent-citation-function

Function for modifying a citation just inserted in the mail buffer. This can also be a list of functions. Each function can find the citation between (point) and (mark t). And each function should leave point and mark around the citation text as modified.

message-signature

String to be inserted at the end of the message buffer. If t (which is the default), the message-signature-file file will be inserted instead. If a function, the result from the function will be used instead. If a form, the result from the form will be used instead. If this variable is nil, no signature will be inserted at all.

message-signature-file

File containing the signature to be inserted at the end of the buffer. The default

          is ‘~/.signature’.

Note that RFC1036bis says that a signature should be preceded by the three characters ‘-- ’ on a line by themselves. This is to make it easier for the recipient to automatically recognize and process the signature. So don’t remove those characters, even though you might feel that they ruin your beautiful design, like, totally.

Also note that no signature should be more than four lines long. Including ASCII graphics is an efficient way to get everybody to believe that you are silly and have nothing important to say.

2.4 Various Commands

C-c C-r Caesar rotate (aka. rot13) the current message (message-caesar-buffer-body). If narrowing is in effect, just rotate the visible portion of the buffer. A numerical prefix says how many places to rotate the text. The default is 13.

C-c C-e Elide the text between point and mark (message-elide-region). The text is killed and an ellipsis (‘[...]’) will be inserted in its place.

C-c C-z Kill all the text up to the signature, or if that’s missing, up to the end of the

          message (message-kill-to-signature).

C-c C-v Delete all text in the body of the message that is outside the region (message-

          delete-not-region).

M-RET Insert four newlines, and then reformat if inside quoted text.

Here’s an example:

               > This is some quoted text.  And here's more quoted text.

If point is before ‘And’ and you press M-RET, you’ll get:

               > This is some quoted text.
               *
               > And here's more quoted text.

‘*’ says where point will be placed.

C-c C-t Insert a To header that contains the Reply-To or From header of the message you’re following up (message-insert-to).

C-c C-n Insert a Newsgroups header that reflects the Followup-To or Newsgroups

header of the article you’re replying to (message-insert-newsgroups).

C-c M-r Rename the buffer (message-rename-buffer). If given a prefix, prompt for a new buffer name.

2.5 Sending

C-c C-c Send the message and bury the current buffer (message-send-and-exit).

C-c C-s Send the message (message-send).

C-c C-d Bury the message buffer and exit (message-dont-send).

C-c C-k Kill the message buffer and exit (message-kill-buffer).

2.6 Mail Aliases

The message-mail-alias-type variable controls what type of mail alias expansion to use. Currently only one form is supported—Message uses mailabbrev to handle mail aliases. If this variable is nil, no mail alias expansion will be performed.

mailabbrev works by parsing the ‘/etc/mailrc’ and ‘~/.mailrc’ files. These files look like:

     alias lmi "Lars Magne Ingebrigtsen <larsi@ifi.uio.no>"
     alias ding "ding@ifi.uio.no (ding mailing list)"

After adding lines like this to your ‘~/.mailrc’ file, you should be able to just write ‘lmi’ in the To or Cc (and so on) headers and press SPC to expand the alias.

No expansion will be performed upon sending of the message—all expansions have to be done explicitly.

3 Variables

3.1 Message Headers

Message is quite aggressive on the message generation front. It has to be – it’s a combined news and mail agent. To be able to send combined messages, it has to generate all headers itself (instead of letting the mail/news system do it) to ensure that mail and news copies of messages look sufficiently similar.

message-generate-headers-first

If non-nil, generate all headers before starting to compose the message.

message-from-style

Specifies how From headers should look. There are four valid values:

nil  Just the address – ‘king@grassland.com’.
parens  ‘king@grassland.com (Elvis Parsley)’.
angles  ‘Elvis Parsley <king@grassland.com>’.

default Look like angles if that doesn’t require quoting, and parens if it does. If even parens requires quoting, use angles anyway.

message-deletable-headers

Headers in this list that were previously generated by Message will be deleted before posting. Let’s say you post an article. Then you decide to post it again to some other group, you naughty boy, so you jump back to the *post-buf* buffer, edit the Newsgroups line, and ship it off again. By default, this variable makes sure that the old generated Message-ID is deleted, and a new one generated. If this isn’t done, the entire empire would probably crumble, anarchy would prevail, and cats would start walking on two legs and rule the world. Allegedly.

message-default-headers

This string is inserted at the end of the headers in all message buffers.

message-subject-re-regexp

Responses to messages have subjects that start with ‘Re: ’. This is not an abbreviation of the English word “response”, but in Latin, and means “in response to”. Some illiterate nincompoops have failed to grasp this fact, and have “internationalized” their software to use abonimations like ‘Aw: ’ (“antwort”) or ‘Sv: ’ (“svar”) instead, which is meaningless and evil. However, you may have to deal with users that use these evil tools, in which case you may set this variable to a regexp that matches these prefixes. Myself, I just throw away non-compliant mail.

3.2 Mail Headers

message-required-mail-headers

See Section 3.4 [News Headers], page 12, for the syntax of this variable.

It is (From Date Subject (optional . In-Reply-To) Message-ID Lines
(optional . X-Mailer)) by default.

message-ignored-mail-headers

Regexp of headers to be removed before mailing. The default is

‘^[GF]cc:\\|^Resent-Fcc:’.
message-default-mail-headers

This string is inserted at the end of the headers in all message buffers that are initialized as mail.

3.3 Mail Variables

message-send-mail-function

Function used to send the current buffer as mail. The default is messagesend-mail-with-sendmail. If you prefer using MH instead, set this variable

to message-send-mail-with-mh.
message-mh-deletable-headers

Most versions of MH doesn’t like being fed messages that contain the headers in this variable. If this variable is non-nil (which is the default), these headers will be removed before mailing when sending messages via MH. Set it to nil if your MH can handle these headers.

3.4 News Headers

message-required-news-headers a list of header symbols. These headers will either be automatically generated, or, if that’s impossible, they will be prompted for. The following symbols are valid:

From This required header will be filled out with the result of the message-make-

from function, which depends on the message-from-style, user-full-name,
user-mail-address variables.

Subject This required header will be prompted for if not present already.

Newsgroups

This required header says which newsgroups the article is to be posted to. If it isn’t present already, it will be prompted for.

Organization

This optional header will be filled out depending on the message-user-

organization variable. message-user-organization-file will be used if

this variable is t. This variable can also be a string (in which case this string will be used), or it can be a function (which will be called with no parameters and should return a string to be used).

Lines This optional header will be computed by Message.

Message-ID

This required header will be generated by Message. A unique ID will be created based on the date, time, user name and system name. Message will use mailhost-address as the fully qualified domain name (FQDN) of the machine if that variable is defined. If not, it will use system-name, which doesn’t report a FQDN on some machines – notably Suns.

X-Newsreader

This optional header will be filled out according to the message-newsreader local variable.

X-Mailer This optional header will be filled out according to the message-mailer local variable, unless there already is an X-Newsreader header present.

In-Reply-To

This optional header is filled out using the Date and From header of the article being replied to.

Expires This extremely optional header will be inserted according to the messageexpires variable. It is highly deprecated and shouldn’t be used unless you know what you’re doing.

Distribution

This optional header is filled out according to the message-distribution-function variable. It is a deprecated and much misunderstood header.

Path This extremely optional header should probably never be used. However, some very old servers require that this header is present. message-user-path further controls how this Path header is to look. If it is nil, use the server name as the leaf node. If it is a string, use the string. If it is neither a string nor nil, use the user name only. However, it is highly unlikely that you should need to fiddle with this variable at all.

In addition, you can enter conses into this list. The car of this cons should be a symbol. This symbol’s name is the name of the header, and the cdr can either be a string to be entered verbatim as the value of this header, or it can be a function to be called. This function should return a string to be inserted. For instance, if you want to insert Mime-Version: 1.0, you should enter (Mime-Version . "1.0") into the list. If you want to insert a funny quote, you could enter something like (X-Yow . yow) into the list. The function yow will then be called without any arguments.

If the list contains a cons where the car of the cons is optional, the cdr of this cons will only be inserted if it is non-nil.

Other variables for customizing outgoing news articles:

message-syntax-checks

Controls what syntax checks should not be performed on outgoing posts. To disable checking of long signatures, for instance, add

     (signature . disabled)

to this list. Valid checks are:

subject-cmsg

Check the subject for commands.

sender Insert a new Sender header if the From header looks odd.

multiple-headers

Check for the existence of multiple equal headers. sendsys Check for the existence of version and sendsys commands.

message-id

Check whether the Message-ID looks ok.

from Check whether the From header seems nice.

long-lines

Check for too long lines.

control-chars

Check for invalid characters.

size Check for excessive size.

new-text Check whether there is any new text in the messages.

signature

Check the length of the signature.

approved Check whether the article has an Approved header, which is something only moderators should include.

empty Check whether the article is empty.

empty-headers

Check whether any of the headers are empty.

existing-newsgroups

Check whether the newsgroups mentioned in the Newsgroups and Followup-To headers exist.

valid-newsgroups

Check whether the Newsgroups and Followup-to headers are valid syntactically.

repeated-newsgroups

Check whether the Newsgroups and Followup-to headers contains repeated group names.

shorten-followup-to

Check whether to add a Followup-to header to shorten the number of groups to post to.

All these conditions are checked by default.

message-ignored-news-headers

Regexp of headers to be removed before posting. The default is

‘^NNTP-Posting-Host:\\|^Xref:\\|^[BGF]cc:\\|^Resent-Fcc:’.
message-default-news-headers

This string is inserted at the end of the headers in all message buffers that are initialized as news.

3.5 News Variables

message-send-news-function

Function used to send the current buffer as news. The default is message-

send-news.
message-post-method

Gnusish select method (see the Gnus manual for details) used for posting a prepared news message.

3.6 Various Message Variables

message-signature-separator

Regexp matching the signature separator. It is ‘^-- *$’ by default.

mail-header-separator

String used to separate the headers from the body. It is ‘--text follows this line--’ by default.

message-directory

Directory used by many mailey things. The default is ‘~/Mail/’.

message-signature-setup-hook

Hook run when initializing the message buffer. It is run after the headers have been inserted but before the signature has been inserted.

message-setup-hook

Hook run as the last thing when the message buffer has been initialized, but before yanked text is inserted.

message-header-setup-hook

Hook called narrowed to the headers after initializing the headers. For instance, if you’re running Gnus and wish to insert a ‘Mail-Copies-To’ header in all your news articles and all messages you send to mailing lists, you could do something like the following:

     (defun my-message-header-setup-hook ()
       (let ((group (or gnus-newsgroup-name "")))
         (when (or (message-fetch-field "newsgroups")
                   (gnus-group-find-parameter group 'to-address)
                   (gnus-group-find-parameter group 'to-list))
           (insert "Mail-Copies-To: never\n"))))
     (add-hook 'message-header-setup-hook
               'my-message-header-setup-hook)
message-send-hook

Hook run before sending messages. If you want to add certain headers before sending, you can use the messageadd-header function in this hook. For instance:

     (add-hook 'message-send-hook 'my-message-add-content)
     (defun my-message-add-content ()
       (message-add-header
        "Mime-Version: 1.0"
        "Content-Type: text/plain"
        "Content-Transfer-Encoding: 7bit"))

This function won’t add the header if the header is already present.

message-send-mail-hook

Hook run before sending mail messages.

message-send-news-hook

Hook run before sending news messages.

message-sent-hook

Hook run after sending messages.

message-mode-syntax-table

Syntax table used in message mode buffers.

message-send-method-alist

Alist of ways to send outgoing messages. Each element has the form

     (TYPE PREDICATE FUNCTION)

type A symbol that names the method.

predicate A function called without any parameters to determine whether the message is a message of type type.

function A function to be called if predicate returns non-nil. function is

called with one parameter – the prefix.

     ((news message-news-p message-send-via-news)
      (mail message-mail-p message-send-via-mail))

3.7 Sending Variables

message-fcc-handler-function

A function called to save outgoing articles. This function will be called with the name of the file to store the article in. The default function is message-output which saves in Unix mailbox format.

message-courtesy-message

When sending combined messages, this string is inserted at the start of the mailed copy. If the string contains the format spec ‘%s’, the newsgroups the article has been posted to will be inserted there. If this variable is nil, no such courtesy message will be added. The default value is ‘"The following

message is a courtesy copy of an article\nthat has been posted to %s
as well.\n\n"’.

3.8 Message Buffers

Message will generate new buffers with unique buffer names when you request a message buffer. When you send the message, the buffer isn’t normally killed off. Its name is changed and a certain number of old message buffers are kept alive.

message-generate-new-buffers

If non-nil, generate new buffers. The default is t. If this is a function, call that function with three parameters: The type, the to address and the group name. (Any of these may be nil.) The function should return the new buffer name.

message-max-buffers

This variable says how many old message buffers to keep. If there are more message buffers than this, the oldest buffer will be killed. The default is 10. If this variable is nil, no old message buffers will ever be killed.

message-send-rename-function

After sending a message, the buffer is renamed from, for instance, ‘*reply to Lars*’ to ‘*sent reply to Lars*’. If you don’t like this, set this variable to a function that renames the buffer in a manner you like. If you don’t want to rename the buffer at all, you can say:

               (setq message-send-rename-function 'ignore)
message-kill-buffer-on-exit

If non-nil, kill the buffer immediately on exit.

3.9 Message Actions

When Message is being used from a news/mail reader, the reader is likely to want to perform some task after the message has been sent. Perhaps return to the previous window configuration or mark an article as replied.

The user may exit from the message buffer in various ways. The most common is C-c C-c, which sends the message and exits. Other possibilities are C-c C-s which just sends the message, C-c C-d which postpones the message editing and buries the message buffer, and C-c C-k which kills the message buffer. Each of these actions have lists associated with them that contains actions to be executed: message-send-actions, message-exit-

actions, message-postpone-actions, and message-kill-actions.

Message provides a function to interface with these lists: message-add-action. The first parameter is the action to be added, and the rest of the arguments are which lists to add this action to. Here’s an example from Gnus:

       (message-add-action
        `(set-window-configuration ,(current-window-configuration))
        'exit 'postpone 'kill)

This restores the Gnus window configuration when the message buffer is killed, postponed or exited.

An action can be either: a normal function, or a list where the car is a function and the cdr is the list of arguments, or a form to be evaled.

18 Message 5.7 Manual

4 Compatibility

Message uses virtually only its own variables—older mail- variables aren’t consulted. To force Message to take those variables into account, you can put the following in your

.emacs file:
     (require 'messcompat)

This will initialize many Message variables from the values in the corresponding mail variables.

20 Message 5.7 Manual

5 Appendices

5.1 Responses

To determine where a message is to go, the following algorithm is used by default. reply A reply is when you want to respond just to the person who sent the message

via mail. There will only be one recipient. To determine who the recipient will be, the following headers are consulted, in turn:

Reply-To
From

wide reply A wide reply is a mail response that includes all entities mentioned in the message you are responded to. All mailboxes from the following headers will be concatenated to form the outgoing To/Cc headers: From (unless there’s a Reply-To, in which case that is used instead).

Cc
To

If a Mail-Copies-To header is present, it will also be included in the list of mailboxes. If this header is ‘never’, that means that the From (or Reply-To) mailbox will be suppressed.

followup A followup is a response sent via news. The following headers (listed in order of precedence) determine where the response is to be sent:

Followup-To
Newsgroups

If a Mail-Copies-To header is present, it will be used as the basis of the new Cc header, except if this header is ‘never’.

22 Message 5.7 Manual Chapter 6: Index 23

6 Index

A

aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 approved . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

C

compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

D

Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

E

Expires . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

F

From . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

L

Lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 long lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

M

mail aliases. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 mail-header-separator . . . . . . . . . . . . . . . . . . . . . . . . 15 mail-host-address . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 message-add-header . . . . . . . . . . . . . . . . . . . . . . . . . . 15 message-bounce. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 message-caesar-buffer-body . . . . . . . . . . . . . . . . . . . . 9 message-cancel-news . . . . . . . . . . . . . . . . . . . . . . . . . . 4 message-citation-line-function . . . . . . . . . . . . . . . . . . 8 message-cite-function. . . . . . . . . . . . . . . . . . . . . . . . . . 8 message-cite-original . . . . . . . . . . . . . . . . . . . . . . . . . . 8 message-cite-original-without-signature . . . . . . . . . 8 message-courtesy-message . . . . . . . . . . . . . . . . . . . . 16 message-default-headers . . . . . . . . . . . . . . . . . . . . . . 11 message-default-mail-headers . . . . . . . . . . . . . . . . . 12 message-default-news-headers . . . . . . . . . . . . . . . . . 14 message-deletable-headers . . . . . . . . . . . . . . . . . . . . 11 message-delete-not-region. . . . . . . . . . . . . . . . . . . . . . 9 message-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 message-dont-send . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 message-elide-region. . . . . . . . . . . . . . . . . . . . . . . . . . . 9 message-exit-actions . . . . . . . . . . . . . . . . . . . . . . . . . 17 message-fcc-handler-function. . . . . . . . . . . . . . . . . . 16 message-fill-yanked-message . . . . . . . . . . . . . . . . . . . 8 message-followup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 message-followup-to-function . . . . . . . . . . . . . . . . . . 4 message-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 message-forward-end-separator . . . . . . . . . . . . . . 4, 5 message-forward-start-separator. . . . . . . . . . . . . . . . 4 message-forward-subject-author-subject. . . . . . . . . 5 message-from-style . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 message-generate-headers-first . . . . . . . . . . . . . . . . 11 message-generate-new-buffers . . . . . . . . . . . . . . . . . 17 message-goto-bcc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 message-goto-body . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 message-goto-cc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 message-goto-distribution. . . . . . . . . . . . . . . . . . . . . . 7 message-goto-fcc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 message-goto-followup-to . . . . . . . . . . . . . . . . . . . . . . 7 message-goto-keywords . . . . . . . . . . . . . . . . . . . . . . . . 7 message-goto-newsgroups . . . . . . . . . . . . . . . . . . . . . . 7 message-goto-reply-to . . . . . . . . . . . . . . . . . . . . . . . . . 7 message-goto-signature . . . . . . . . . . . . . . . . . . . . . . . . 7 message-goto-subject . . . . . . . . . . . . . . . . . . . . . . . . . . 7 message-goto-summary . . . . . . . . . . . . . . . . . . . . . . . . 7 message-goto-to. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 message-header-setup-hook . . . . . . . . . . . . . . . . . . . 15 Message-ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 message-ignored-bounced-headers . . . . . . . . . . . . . . 5 message-ignored-cited-headers . . . . . . . . . . . . . . . . . 8 message-ignored-mail-headers . . . . . . . . . . . . . . . . . 12 message-ignored-news-headers . . . . . . . . . . . . . . . . 14 message-ignored-resent-headers . . . . . . . . . . . . . . . . 5 message-ignored-supersedes-headers . . . . . . . . . . . . 4 message-included-forward-headers . . . . . . . . . . . . . . 5 message-indent-citation-function . . . . . . . . . . . . . . . 8 message-indentation-spaces . . . . . . . . . . . . . . . . . . . . 8 message-insert-newsgroups. . . . . . . . . . . . . . . . . . . . . 9 message-insert-signature . . . . . . . . . . . . . . . . . . . . . . . 8 message-insert-to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 message-kill-actions . . . . . . . . . . . . . . . . . . . . . . . . . . 17 message-kill-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 message-kill-buffer-on-exit . . . . . . . . . . . . . . . . . . . . 17 message-kill-to-signature . . . . . . . . . . . . . . . . . . . . . . 9 message-mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 message-mail-alias-type . . . . . . . . . . . . . . . . . . . . . . 10 message-make-forward-subject-function . . . . . . . . . 5 message-max-buffers . . . . . . . . . . . . . . . . . . . . . . . . . 17 message-mh-deletable-headers. . . . . . . . . . . . . . . . . 12 message-mode-syntax-table . . . . . . . . . . . . . . . . . . . 16

24 Message 5.7 Manual

message-news. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 message-post-method . . . . . . . . . . . . . . . . . . . . . . . . 15 message-postpone-actions. . . . . . . . . . . . . . . . . . . . . 17 message-rename-buffer . . . . . . . . . . . . . . . . . . . . . . . . 9 message-reply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 message-reply-to-function. . . . . . . . . . . . . . . . . . . . . . 3 message-required-mail-headers . . . . . . . . . . . . . . . . 11 message-required-news-headers. . . . . . . . . . . . . . . . 12 message-resend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 message-send . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 message-send-actions . . . . . . . . . . . . . . . . . . . . . . . . . 17 message-send-and-exit. . . . . . . . . . . . . . . . . . . . . . . . 10 message-send-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 message-send-mail-function . . . . . . . . . . . . . . . . . . . 12 message-send-mail-hook . . . . . . . . . . . . . . . . . . . . . . 16 message-send-method-alist. . . . . . . . . . . . . . . . . . . . 16 message-send-news-function . . . . . . . . . . . . . . . . . . 15 message-send-news-hook. . . . . . . . . . . . . . . . . . . . . . 16 message-send-rename-function . . . . . . . . . . . . . . . . 17 message-sent-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 message-setup-hook . . . . . . . . . . . . . . . . . . . . . . . . . . 15 message-signature. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 message-signature-before-forwarded-message …. 5 message-signature-file . . . . . . . . . . . . . . . . . . . . . . . . . 9 message-signature-separator . . . . . . . . . . . . . . . . . . 15 message-signature-setup-hook . . . . . . . . . . . . . . . . . 15 message-subject-re-regexp . . . . . . . . . . . . . . . . . . . . 11 message-supersede . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 message-syntax-checks . . . . . . . . . . . . . . . . . . . . . . . 13 message-use-followup-to . . . . . . . . . . . . . . . . . . . . . . . 4 message-wash-forwarded-subjects. . . . . . . . . . . . . . . 5 message-wide-reply. . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 message-wide-reply-to-function. . . . . . . . . . . . . . . . . 4 message-yank-original . . . . . . . . . . . . . . . . . . . . . . . . . 8 message-yank-prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Mime-Version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

N

Newsgroups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

O

organization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

P

path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

Q

quoting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

R

rmail-dont-reply-to-names . . . . . . . . . . . . . . . . . . . . . 4

S

sc-cite-original . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Sender . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 sendsys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Subject . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Sun. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Supercite. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 system-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

U

user-full-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 user-mail-address . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

X

X-Newsreader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

Y

yanking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 yow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

Chapter 7: Key Index 25

7 Key Index

C

C-c ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 C-c C-d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 C-c C-e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 C-c C-f C-b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-f C-c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-f C-d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-f C-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-f C-k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-f C-n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-f C-o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-f C-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-f C-s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-f C-t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-f C-u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 C-c C-k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 C-c C-n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 C-c C-q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 C-c C-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 C-c C-s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 C-c C-t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 C-c C-v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 C-c C-w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 C-c C-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 C-c C-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 C-c M-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

M

M-RET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

                                                message-newline-and-reformat . . . . . . . . . . . . . . . 9

26 Message 5.7 Manual

Short Contents

Message. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2 Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3 Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4 Compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 5 Appendices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 6 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 7 Key Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ii Message 5.7 Manual

Table of Contents

Message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.1 New Mail Message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 New News Message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Reply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.4 Wide Reply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.5 Followup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.6 Canceling News . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.7 Superseding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.8 Forwarding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.9 Resending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.10 Bouncing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

2 Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.1 Header Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2 Movement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.3 Insertion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.4 Various Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.5 Sending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.6 Mail Aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

3 Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

3.1 Message Headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.2 Mail Headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.3 Mail Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.4 News Headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.5 News Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.6 Various Message Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.7 Sending Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.8 Message Buffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.9 Message Actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

4 Compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

5 Appendices. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

5.1 Responses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

6 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

7 Key Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

iv Message 5.7 Manual